OpenStudy (anonymous):

A circle is represented by the equation shown below. (x + 3)2 + (x - 2)2 = 4

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Answer: C. The circle is centered at (-3, 2) and has a radius of 2. how would I find this?

OpenStudy (anonymous):

It is circle with center at (-3,2) and radius 2

OpenStudy (anonymous):

yes but how would I work it out?

OpenStudy (anonymous):

In general equation of circle with center x0 y0 and radius r is (x-x0)^2+(y-y0)^2=r^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the left let x+3=0->x=-3 y-2=0->y=2 the right 4=2^2